can potassium carbonate remove calcium hardness from water instead of Sodium Carbonate,
I have a KCl solution 20g/l with Calcium 500mg/l, I would normally use Na2CO3 to remove the calcium
1
Expert's answer
2020-04-02T09:01:36-0400

can potassium carbonate remove calcium hardness from water instead of Sodium Carbonate,

I have a KCl solution 20g/l with Calcium 500mg/l, I would normally use Na2CO3 to remove the calcium


Solution:

yes you can


now as equation goes

2KCl+CaSO4+H2O=K2SO4(aq)+CaCl2


hence 1 mole of K2CO3 reacts with one mole of CaSO4

20g KCl=20/74=0.27 mol

0.5 g CaSO4=0.5/136= 0.00367 mol